GROUND-STATE SPLITTINGS OF ACTINIDE IONS IN CaF_{2}

\underline{SCOTT R. BROZELL} and RUSSELL M. PITZER, Department of Chemical Physics, The Ohio State University, Columbus, OH 43210.

The energy-level splittings of Am^{2+}, Cm^{3+}, Es^{2+}, Pu^{3+}, and U^{4+} in an octahedral site of CaF_{2} were studied using restricted Hartree-Fock and spin-orbit configuration-interaction calculations. The CaF_{2} host was modeled with a large finite cluster of ions which approximate the Madelung potential of the crystal lattice^{a}. The actinide dopants were treated with relativistic effective core potentials and Gaussian correlation consistent double-zeta plus polarization basis sets. The primary properties of interest are energy splittings and magnetic moments.

aN. W. Winter, M. Ross and R. M. Pitzer J. Phys. Chem. 94, 1172 (1990).